he following table represents fasting blood glucose levels measured in monozygotic twins who were raised apart. Each row represents a separate pair of twins.

Set 1 95 97
Set 2 93 90
Set 3 89 92
Set 4 101 90
Set 5 92 105

What is the covariance of fasting blood glucose between sets of twins?
Answer #1

We need to compute the covariance, which is computed by first computing cross products of the sample data.

Now, with the provided sample data, we need to construct the following table, which will be used for the calculation of the covariance coefficient:

Obs. X Y X; Y
1 95 97 9215
2 93 90 8370
3 89 92 8188
4 101 90 9090
5 92 105 9660
Sum = 470 474 44523

Based on the table above, we compute the following sum of cross-products that will be used in the calculation of the covariance:

$$y = $x* ++ (*) (* = 44523 - x 222780 = -33

Now, the covariance between X and Y is computed using the following expression:

cou(X,Y)= SSXY n - 1 = 5-1 = -8.25

Therefore, based on the sample data provided, it is found that the sample covariance coefficient is

cov(X,Y) = -8.25

Let me know in the comments if anything is not clear. I will reply ASAP! Please do upvote if satisfied!